if f(x)=x^2, g(x)=Sinx and h(x)=x/2 then f(g(h(x)))=??

Grade:upto college level

2 Answers

Radhika Batra
10 years ago
f(g(h(x)))
=f(g(x/2))
=f(sin(x/2))
=sin^2(x/2)
jayaram
10 years ago

f(x)=x^2

g(x)=sinx

h(x)=x/2

f(f(h(x)))

a)f(g[x/2]

f[sin x/2]

=(sinx/2)^2

=sin^2 x/2